OpenStudy (anonymous):

Solve for x: log^(5x+9)(x^(2)+6x+9) + log^(x+3) (5x^(2)+24x+27)=4

OpenStudy (anonymous):

\[\log_{5x+9}(x^{2}+6x+9)+\log_{x+3} (5x ^{2}+24x+27)=4 \]

OpenStudy (day_2010):

Use the product rule to get log_bM+log_bN for the first half right? .....so that would be log(5x+9)+log(x^2) ... right? Im new at this

OpenStudy (anonymous):

I think so i'm not sure

OpenStudy (day_2010):

yeah cause if you simplify too soon it will be a mess lol

OpenStudy (anonymous):

so i would set m=log(5x+9) n=log(x+3) for example?

OpenStudy (day_2010):

the (5x+9)(x^2) is being multiplied together right? just clarifying

OpenStudy (anonymous):

ya

OpenStudy (day_2010):

okay, neato! ... so, yes start with the product rule to